OpenStudy (anonymous):

How can i solve integral of negative exponents?? like u^-2.. ? Thank you..

ganeshie8 (ganeshie8):

power rule works for both negative and positive exponents (-1 is the only exception)

ganeshie8 (ganeshie8):

\[\large \int x^n dx = \dfrac{x^{n+1}}{n+1} + C\]

OpenStudy (anonymous):

What if the exponent is -1??

ganeshie8 (ganeshie8):

haha then everything falls apart and the world comes to end

ganeshie8 (ganeshie8):

\[\large \int x^{-1}dx = \int \dfrac{1}{x}dx = \ln |x| + C\]

OpenStudy (larseighner):

\[ \large \int {1 \over x} dx = ln |x| \]

OpenStudy (anonymous):

Ok. Thank you for the answers!! :) :) :)

OpenStudy (anonymous):

Can you give me techniques on how to analyze equations to solve b4 solving them? I'm currently taking up differential equations and I'm having difficulties in integrals..

OpenStudy (larseighner):

The one I alway get wrong: \[ \large \int k dx = kx \] I alway try to differentiate the constant instead of integrating.

ganeshie8 (ganeshie8):

familiar with separation of variables ?

OpenStudy (anonymous):

Yes. But after separating variables, here comes integration and as you said earlier, the world comes to an end..

ganeshie8 (ganeshie8):

Oh yes, integration is a different story - to my knowledge, the most difficult part is solving the DE... not integration

ganeshie8 (ganeshie8):

solving the DE is same as isolating the dependent variable "y"

ganeshie8 (ganeshie8):

usually, in a DE course, you will be given equations such that they turn out into simpler integrals... the main focus is on solving the DE, not on integration.

ganeshie8 (ganeshie8):

knowing below integral tricks are sufficient most of the time : 1) u substitution 2) trig substitution 3) partial fractions `4) integration by parts ` (no need to master this)

OpenStudy (anonymous):

OoooKKK. Guess I'll have to review those topics overnight.. Thank you so so much.. :)

OpenStudy (larseighner):

There is also what Prof, Jerison calls "advanced guessing" \[ \large \int xe^{-x^2} dx \] You can guess \( e^{-x^2} \) and then differentiate to see what needs adjusting. In this case \[\large {d \over {dx}} e^{-x^2} = -2xe^{-x^2}\] and \(-{! \over 2} \) of that is our integrand. So \[ \large \int xe^{-x^2} dx = - {1 \over 2} e^{-x^2} \]

ganeshie8 (ganeshie8):

you're good to start DE course if you know how to take below integral using u substitution : \[\large \int 2xe^{x^2}~ dx\]

OpenStudy (ikram002p):

mmm

ganeshie8 (ganeshie8):

you took Linear algebra before ? @kai021

OpenStudy (anonymous):

Advance algebra, it is.. 1st semester. I think my biggest problem is recalling what previous lessons..

ganeshie8 (ganeshie8):

Don't wry, DE is not that hard... you can review everything parallel

OpenStudy (ikram002p):

how linear is related to this ?

ganeshie8 (ganeshie8):

solving systems of equations, matrix exponentials, matrix methods for inhomogeneous systems... all these require some background in LA

OpenStudy (anonymous):

try. :) i let \[u=e ^{x ^{2}}\] and du=\[=xe ^{x ^{2}}\]dx

ganeshie8 (ganeshie8):

that covers almost 1/4th of any DE course

OpenStudy (ikram002p):

xD

ganeshie8 (ganeshie8):

Wow ! thats a very unusual try !!! but it works xD it never occurred to me before !

ganeshie8 (ganeshie8):

\[\large \int 2xe^{x^2}~ dx \]

ganeshie8 (ganeshie8):

so, you want to let \(\large u = e^{x^2}\) ? \(\large \implies du = e^{x^2} (2x) dx \) \(\large \implies du = 2xe^{x^2}dx \) So, the integral becomes : \[\large \int du\]

OpenStudy (anonymous):

Wait. my approach is wrong.. :3

ganeshie8 (ganeshie8):

which is same as : \[\large \int 1 du\] \[\large u + C\] \[\large e^{x^2} + C\]

ganeshie8 (ganeshie8):

nope, your approach is perfectly correct ! and i must say its the best substitution !

OpenStudy (anonymous):

Should it be u=e^(x^2) and v= x ???

ganeshie8 (ganeshie8):

avoid integration by parts when u can

ganeshie8 (ganeshie8):

always try substitution first; try integration by parts only after exhausting all other options.

OpenStudy (anonymous):

Oh. ok. :) Hope I won't get stuck with integrating tomorrow. We'll have our preliminary exams.. :) Thank you!! :)

ganeshie8 (ganeshie8):

good luck !

OpenStudy (anonymous):

:)

OpenStudy (kainui):

A fairly good way to reason out why x^-1 doesn't work like all the rest is because if you differentiate \[\Large \frac{d}{dx}(x^0)=0*x^{-1}\] Of course, anything to the zero power is a constant, so that's why you get zero here. The problem is we really don't have any standard way to get there haha.

OpenStudy (ikram002p):

ok lol within week ill master DEs :P both ordinary and partial xD